OpenStudy (anonymous):

How would I write this: (3a^2b^3c^5)/(8x^4y^3z) (it's fraction form) with no denominator?

OpenStudy (anonymous):

\[\huge {x \over y^n} \equiv x*y^{-n}\] where n is the power of y.

OpenStudy (anonymous):

\[ (3a^2b^3c^5) \times 2^{-3} \times 7^{-3} \times z{^-1}\]